OpenStudy (anonymous):

PLEASE HELP I need three points for -y=x+0 I HAVE NO IDEA HOW TO FIND THEM

OpenStudy (anonymous):

idk either

OpenStudy (anonymous):

please help lol someone thank you though @CandyCoatedRainDrops

OpenStudy (texaschic101):

-y = x + 0 so pick any number for y....lets pick 3...now we sub -3 = x + 0 -3 = x so if y = 3, the x = -3.....(-3,3) now lets let y be 2....now we sub -y = x + 0 -2 = x + 0 -2 = x so if y = 2, then x = -2 can you find the other set of points ?